# CMRAG: Co-modality–based document retrieval and visual question answering

### v1

@@ Abstract
Retrieval-Augmented Generation (RAG) has become a core paradigm in document question answering tasks. However, existing methods have limitations when dealing with multimodal documents: one category of methods relies on layout analysis and text extraction, which can only utilize explicit text information and struggle to capture images or unstructured content; the other category treats document segmentation as visual input and directly passes it to visual language models (VLMs) for processing, yet it ignores the semantic advantages of text, leading to suboptimal generation results. This paper proposes co-modality–based RAG (CMRAG), which can simultaneously leverage text and images for efficient retrieval and generation. Specifically, we first perform structured parsing on documents to obtain co-modality representations of text segments and image regions. Subsequently, in response to user queries, we retrieve candidate evidence from text and image channels, respectively, and aggregate the results at the cross-modal retrieval level. Finally, we prompt the VLM to generate the final response based on the co-modality retrieval results. Experiments demonstrate that our method significantly outperforms pure-vision–based RAG in visual document question answering tasks. The findings of this paper show that integrating co-modality information into the RAG framework in a unified manner is an effective approach to improving the performance of complex document visual question-answering (VQA) systems.
@@ 1 Introduction
One line of approaches is text-based RAG, which typically relies on layout parsing and text extraction (Xu et al., 2020; Dong et al., 2025; Yang et al., 2025b; Perez & Vizcaino, 2024). These methods first detect document layouts and then extract textual information for subsequent retrieval and generation. While stable at the semantic level, they struggle to handle content such as images and tables. Recently, VLMs (Ghosh et al., 2024; Radford et al., 2021; Alayrac et al., 2022) enable RAG systems to process documents directly as images (Faysse et al., 2025; Yu et al., 2025; Qi et al., 2024a; Wang et al., 2025b), giving rise to vision-based RAG (Huang et al., 2022; Kim et al., 2022; Yu et al., 2024). Specifically, these methods divide document pages into image segments and perform retrieval and reasoning through visual understanding models. Although they capture non-textual information, they often overlook the precise information carried by text, leading to performance bottlenecks.
To overcome these limitations, we propose a novel co-modality–based RAG framework, which unifies text and image modalities, as illustrated in Fig. 1. In this framework, we first parse documents to extract structured text and image segments. For a given query, we then perform retrieval in both text and visual spaces, ensuring that semantic matching from text and perceptual grounding from images are simultaneously leveraged. Finally, we feed the co-modality evidence into a VLM to integrate information and generate answers. In this way, our method not only handles cross-modal reasoning that single-modal RAG cannot cover but also achieves significant performance improvements on standard visual document question answering benchmarks.
The main contributions of this paper are summarized as follows:
  • We propose a novel RAG framework that jointly leverages text and image modalities to enhance performance on complex visual document QA tasks.
  • We design a co-modality–based retrieval mechanism that enables complementary use of text and image evidence during retrieval, leading to more comprehensive knowledge grounding.
  • We conduct extensive experiments on document VQA benchmarks, and the results demonstrate that our approach significantly outperforms existing single-modal RAG methods in both accuracy and robustness.
@@ 3 Methodology
In this section, we introduce the problem definition of document–based VQA and our proposed CMRAG framework in detail.
@@ 3.1 Problem definition
In a document–based VQA system, a collection of visual documents provides evidence for answering given queries. Formally, all pages are treated as candidate evidence $\mathcal{D}=\{p_{1},p_{2},\dots,p_{M}\}$, where $M$ denotes the total number of pages across the document collection. For a given query $q$, the retriever $\mathcal{R}$ identifies the top-$k$ relevant pages $P_{k}=\{p_{1},p_{2},\dots,p_{k}\}$ according to a similarity score $s(q,p_{i})$ ($\forall p_{i}\in P_{k}$). Once the top-$k$ evidence pages are retrieved, they are combined with the query into a prompt $\mathcal{P}(q,P_{k})$, which is then passed to a generator model $\mathcal{G}$, often instantiated as a vision–language model (VLM). The generator integrates the multimodal evidence and produces the final answer $\hat{a}=\mathcal{G}(\mathcal{P}(q,P_{k}))$.
3.2 VLM–based document parsing
Advanced vision–language models (VLMs) can directly parse document pages, simultaneously detecting layouts, extracting text, and identifying localized visual regions. For each page $p_{i}$, the parsing process produces three complementary components: the entire image $I_{i}$, which preserves global layout and contextual relationships; the parsed sub-images $SI_{i}$, which capture localized elements such as figures, tables, and diagrams; and the parsed text $T_{i}$, which provides precise semantic grounding for fine-grained reasoning. Thus, each page can be represented as $p_{i}=\{I_{i},SI_{i},T_{i}\}$, ensuring that both global structure and detailed content are available for retrieval and downstream answer generation.
3.3 Co-modality late interaction
Late interaction. Given a query $q$ and document page image $p$, the encoder $\mathcal{E}$ project them into a high-dimension space $\mathbb{R}^{D}$, i.e., $\mathbf{q}=\mathcal{E}(q)\in\mathbb{R}^{N_{q}\times D}$ and $\mathbf{p}=\mathcal{E}(p)\in\mathbb{R}^{N_{p}\times D}$, where $N_{q}$ and $N_{p}$ denote the number of vectors in the query and document embeddings, respectively. Hence, the late interaction (LI) (Khattab & Zaharia, 2020; Santhanam et al., 2022; Clavié, 2025; Faysse et al., 2025) can be calculated as follows:
$\textbf{LI}(q,p)=\sum_{i=1}^{N_{q}}\max_{j\in\{1,2,...,N_{p}\}}\langle\mathbf{q}^{i},\mathbf{p}^{j}\rangle,$ (1)
where $\langle\cdot,\cdot\rangle$ denote dot production, and $\mathbf{q}^{i}$ and $\mathbf{p}^{j}$ are $i^{\text{th}}$ and $j^{\text{th}}$ vector in the query and page embeddings, respectively. Compared with the mean pooling method (Karpukhin et al., 2020; Yu et al., 2025), LI can consider token-level similarity for detailed matching (Wan et al., 2025).
Co-modality late interaction. Furthermore, inspired by CLAMR (Wan et al., 2025), we consider retrieving evidence based on similarity between the query $q$ and the co-modality document content $I$ and $T$. Formally, we encode the entire page image $I$ and extracted text $T$ as $\mathbf{p}^{I}=\mathcal{E}(I)$ and $\mathbf{p}^{T}=\mathcal{E}(T)$, respectively. Therefore, we adopt co-modality late interaction (CoLI) to calculate the similarity, as follows:
$\textbf{CoLI}(q,p)=\max_{m\in\{I,T\}}\sum_{i=1}^{N_{q}}\max_{j\in\{1,2,...,N_{p}\}}\langle\mathbf{q}^{i},\mathbf{p}^{m,j}\rangle,$ (2)
where $\mathbf{p}^{m,j}$ is $j^{\text{th}}$ vector in the $m^{\text{th}}$ modality document embedding.
Contrastive learning. Let $\{(q_{i},p_{i})\}_{i=1}^{b}$ denote a data batch where each query $q_{i}$ ($\forall i\in\{1,2,...,b\}$) corresponds to a document page $p_{i}$ and $b$ denotes the batch size. The loss (Oord et al., 2018) can be represented as follows:
$\mathcal{L}=-\frac{1}{b}\sum_{i=1}^{b}\log\frac{\exp(s_{i,i}/\tau)}{\sum_{j=1}^{b}\exp(s_{i,j}/\tau)},$ (3)
where $s_{i,j}$ ($\forall i,j\in\{1,2,...,b\}$) denote the similarity score between $i^{\text{th}}$ query and $j^{\text{th}}$ document page and $\tau$ the temperature.
3.4 Co-modality–based generation
Recent advances in VLMs (Bai et al., 2025; Du et al., 2025) enable them to directly process naive-resolution page images without the need for splitting each page into smaller sub-images, which often risks losing global context. In addition, modern VLMs can accept multiple images simultaneously, meaning that all retrieved pages can be jointly fed into the model. This design facilitates cross-page reasoning, where evidence from different parts of the document can be integrated to produce more accurate and comprehensive answers. Therefore, we combine the retrieved images and text into the prompt and leverage the VLM to generate the final answer based on the co-modality input.
@@ 4.1 Experiment setup
Datasets. We consider two datasets: MMLongBench (Ma et al., 2024b) and SlideVQA (Tanaka et al., 2023). These datasets comprise various query types (multi-hop and single-hop) and answer types (multi-span and single-span), which can be used to evaluate the capacities of VLMs. The statistics of the datasets are summarized in Tab. 1. We follow the implementation of (Wang et al., 2025a) and use refined queries of SlideVQA.
Method #Files #Images #QAs #Images/Q
MMLongBench 135 6,529 1,082 1.00
SlideVQA 400 8,000 2,020 1.26
Table 1: Statistics of evaluation datasets.
Baselines. To evaluate the proposed method, we first consider three evaluation scenarios: (1) Oracle (Entire images), (2) Oracle (Sub-images + parsed text), and (3) Oracle (Entire images + parsed text). We consider using the ground truth document pages, with the entire images, parsed sub-images plus text, and entire images plus parsed text as the input of the VLM, respectively.
LLM-as-a-Judge. The gold answers in the benchmark datasets consist of multiple spans or exhibit diverse valid expressions. Traditional evaluation metrics, such as Exact Match (EM) and token-level F1, struggle to capture the semantic equivalence between generated responses and these multi-span gold answers. As a result, these metrics may underestimate the actual performance of models. To address this limitation, we employ a large language model (LLM) as an automatic judge to assess whether a generated answer is correct, providing a more flexible and reliable evaluation framework, enabling fairer comparisons across different models.
Implementation details. We implemented our CMRAG framework using Qwen2.5-VL-7B-Instruct Bai et al. (2025) as the backbone VLM for parsing documents and answering queries. In addition, we adopted Qwen2.5-7B-Instruct (Yang et al., 2025a) as the judge. To ensure deterministic outputs and eliminate variability due to random sampling, we set the temperature to 0.0 during decoding. We present the prompts used in this study in the appendix.
Method MMLongBench SlideVQA
Oracle (Entire images) 40.39 81.58
Oracle (Sub-images + parsed text) 29.11 58.76
Oracle (Entire images + parsed text) 46.86 81.68
Table 2: Main results. Bold values represent the best scores.

### v2

# CMRAG: Co-modality–based visual document retrieval and question answering

@@ Abstract
Retrieval-Augmented Generation (RAG) has become a core paradigm in document question answering tasks. However, existing methods have limitations when dealing with multimodal documents: one category of methods relies on layout analysis and text extraction, which can only utilize explicit text information and struggle to capture images or unstructured content; the other category treats document segmentation as visual input and directly passes it to visual language models (VLMs) for processing, yet it ignores the semantic advantages of text, leading to suboptimal retrieval and generation results. To address these research gaps, we propose Co-Modality–based RAG (CMRAG) framework, which can simultaneously leverage texts and images for more accurate retrieval and generation. Our framework includes two key components: (1) a Unified Encoding Model (UEM) that projects queries, parsed text, and images into a shared embedding space via triplet-based training, and (2) a Unified Co-Modality–informed Retrieval (UCMR) method that statistically normalizes similarity scores to effectively fuse cross-modal signals. To support research in this direction, we further construct and release a large-scale triplet dataset of (query, text, image) examples. Experiments demonstrate that our proposed framework consistently outperforms single-modality–based RAG in multiple visual document question-answering (VDQA) benchmarks. The findings of this paper show that integrating co-modality information into the RAG framework in a unified manner is an effective approach to improving the performance of complex VDQA systems.
@@ 1 Introduction
One line of approaches is text-based RAG, which typically relies on layout parsing and text extraction (Xu et al., 2020; Dong et al., 2025; Yang et al., 2025b; Perez & Vizcaino, 2024). These methods first detect document layouts and then extract textual information for subsequent retrieval and generation, as shown in Fig. 1(a). While stable at the semantic level, they struggle to handle content such as images and tables. Recently, VLMs (Ghosh et al., 2024; Radford et al., 2021; Alayrac et al., 2022) enable RAG systems to process documents directly as images (Faysse et al., 2025; Yu et al., 2025; Qi et al., 2024a; Wang et al., 2025b), giving rise to vision-based RAG (Huang et al., 2022; Kim et al., 2022; Yu et al., 2025). Specifically, as shown in Fig. 1(b), these methods divide document pages into image segments and perform retrieval and reasoning through visual understanding models. Although they capture non-textual information, they often overlook the precise information carried by text, leading to performance bottlenecks.
To overcome these limitations, we propose a novel co-modality–based RAG (CMRAG) framework, which unifies text and image modalities, as illustrated in Fig. 1(c). In this framework, we first parse documents to extract structured text and image segments. For a given query, we propose a novel retrieval pipeline named CMRAG-Retrieval (CMRAG-R) to retrieve both text and visual representations, ensuring that semantic matching from text and perceptual grounding from images are simultaneously leveraged. Finally, we feed the co-modality evidence into a VLM to integrate information and generate answers.
The CMRAG-R consists of two key components: (1) A unified encoding model (UEM) that projects queries, images, and parsed texts into a shared, comparable latent space. This model is trained using a triplet-based objective with a sigmoid loss to ensure robust alignment across modalities. (2) A unified co-modality-informed retrieval method (UCMR) that normalizes similarity scores to mitigate distributional discrepancies between modalities, enabling a more effective retrieval process. The main contributions of this paper are summarized as follows:
  • We propose CMRAG, a novel co-modality–based RAG framework that leverages both text and image representations for significantly improved retrieval and generation on visual documents.
  • We propose UEM, a unified encoding model that uses a single set of encoders for all modalities, which is trained end-to-end using a pairwise sigmoid loss on query-text and query-image triplets to create a unified embedding space.
  • We propose UCMR, a unified co-modality-informed retrieval method that employs statistical normalization to effectively combine visual and textual similarity scores, addressing the inherent challenges of cross-modal score fusion.
  • We parse and release a large-scale triple dataset constructed from an open-source visual document corpus, providing (query, image, text) triplets to facilitate future research in co-modality learning for the community.
  • We conduct extensive experiments on multiple VDQA benchmarks, showing that our method consistently outperforms strong single-modality RAG baselines and demonstrates the effectiveness of co-modality integration.
@@ 3 Methodology
In this section, we introduce the problem definition, the proposed CMRAG framework, the design and training method of the CMRAG-retrieval, and computational cost analysis.
@@ 3.1 Problem definition
We formulate the task of Visual Document Question Answering (VDQA) within a Retrieval–Augmented Generation (RAG) framework. A collection of visual documents (e.g., PDFs, scanned articles) serves as the knowledge source for answering user queries. As illustrated in Fig. 2(a), each document page $p_{i}$ is first parsed by a Vision Language Model (VLM) $\mathcal{V}$ to extract its structured multimodal content. This parsing step produces both a visual representation $I_{i}$ (which represents the entire page image in this study) and a textual representation $T_{i}$, such that ${I_{i},T_{i}}=\mathcal{V}(p_{i})$. The complete set of all parsed pages constitutes the candidate evidence pool, denoted as $\mathcal{D}=\{\{I_{1},T_{1}\},\{I_{2},T_{2}\},\dots,\{I_{M},T_{M}\}\}$, where $M$ is the total number of pages. For a given query $q$, the objective of the retriever $\mathcal{R}$ is to identify the top-$k$ most relevant pages $P_{k}$ based on a multimodal similarity score $s(q,\{I_{i},T_{i}\})$.
Considering recent advances in VLMs (Bai et al., 2025; Du et al., 2025) enable them to directly process multiple naive-resolution page images, the retrieved evidence $P_{k}$ is directly combined with the query $q$ into a structured prompt $\mathcal{P}(q,P_{k})$, which is then fed into a generator model $\mathcal{G}$, typically a powerful VLM, to produce the final answer: $\hat{a}=\mathcal{G}(\mathcal{P}(q,P_{k}))$ (as shown in Fig. 2(d)). A comprehensive list of notations is provided in Appendix A. Also, we introduce all the prompt templates used in this study in Appendix B.
Figure 2: An overview of the proposed CMRAG framework. (a) A VLM is prompted to parse visual documents offline. (b) Images, parsed texts, and given queries are encoded uniformly in a shared space. Images and texts can be encoded and indexed offline to accelerate the online RAG systems. (c) The calculated similarity scores of visual and textual modalities are unified to a comparable distribution, ensuring a more accurate retrieval. (d) A VLM generator is prompted to generate the final answer based on the query and retrieved evidence.
3.2 CMRAG-Retrieval
In this section, we present a unified model architecture for VDQA, where parsed texts are long and semantically diverse. Existing image–text pretraining models, such as CLIP (Radford et al., 2021), SigLIP (Zhai et al., 2023) and SigLIP2 (Tschannen et al., 2025), are optimized for short texts and degrade on long documents. In addition, they remain biased toward strong image–text correlations in pretraining data. In VDQA, such correlations are sparse, as each image relates to only a small portion of the document. To address this, we propose the CMRAG-R, which jointly embeds queries, images, and parsed texts into a shared latent space for efficient document-level retrieval.
Unified encoding model (UEM). A streamlined yet expressive framework that jointly embeds queries, parsed document texts, and document images into a shared latent space, substantially simplifying the retrieval process. Built on the SigLIP backbone (Zhai et al., 2023), our design integrates three encoders—query $\mathcal{E}^{q}$, image $\mathcal{E}^{I}$, and text $\mathcal{E}^{T}$—into a unified architecture (Fig. 2(b)). We directly reuse the pretrained $\mathcal{E}^{q}$ and $\mathcal{E}^{I}$ from SigLIP to preserve the powerful multimodal alignment acquired at scale, and initialize $\mathcal{E}^{T}$ as a length-extended copy of $\mathcal{E}^{q}$ to accommodate long parsed texts. This yields query, image, and text embeddings:
$\mathbf{q}=\mathcal{E}^{q}(q)\in\mathbb{R}^{1\times d},\quad\mathbf{I}=\mathcal{E}^{I}(I)\in\mathbb{R}^{1\times d},\quad\mathbf{T}=\mathcal{E}^{T}(T)\in\mathbb{R}^{1\times d},$ (1)
where $d$ denotes the embedding dimension. A key benefit of UEM is that a single forward pass of $\mathcal{E}^{q}$ suffices to produce a query representation that can be directly compared against both $\mathbf{T}$ and $\mathbf{I}$, enabling efficient cross-modal retrieval without additional computational overhead (detailed in Section 3.3).
To train UEM and align the three modalities within a shared latent space, we introduce the Dual-Sigmoid Alignment (DSA) loss, a pairwise formulation based on the sigmoid contrastive objective. Given a batch of $b$ triplets $\{(q_{i},I_{i},T_{i})\}_{i=1}^{b}$, the text-query loss is:
$\mathcal{L}^{T}=-\frac{1}{b}\sum_{i=1}^{b}\sum_{j=1}^{b}\log\frac{1}{1+\exp\{\gamma_{ij}(-\tau z^{T}_{ij}+\eta)\}},$ (2)
where $z^{T}_{ij}=\langle\mathbf{q}_{i},\mathbf{T}_{j}\rangle$ is the inner product between the $i^{\text{th}}$ query and $j^{\text{th}}$ text embedding, $\gamma_{ij}\in\{+1,-1\}$ indicates whether the pair is positive or negative, and $\tau,\eta$ are learnable temperature and bias. An analogous loss $\mathcal{L}^{I}$ is defined for query–image pairs with $z^{I}_{ij}=\langle\mathbf{q}_{i},\mathbf{I}_{j}\rangle$. Importantly, we update only $\mathcal{E}^{T}$ during training, keeping $\mathcal{E}^{q}$ and $\mathcal{E}^{I}$ frozen to preserve their pretrained multimodal alignment. A symmetric contrastive regularization further encourages $\mathcal{E}^{T}$ to stay consistent with the frozen encoders, yielding a unified embedding space. The final objective is:
$\mathcal{L}=\lambda\mathcal{L}^{T}+(1-\lambda)\mathcal{L}^{I},$ (3)
where $\lambda\in[0,1]$ controls the relative weight of text and image alignment. Full training details are reported in Appendix D.
Unified co-modality–informed retrieval (UCMR). The core of our framework lies in the retrieval mechanism that effectively leverages the co-modality representations. As shown in Fig. 2(b), given a user query $q$, we first encode it into a dense vector representation $\mathbf{q}$. Offline, the visual and textual components of each document page $p_{i}$ are independently encoded into vectors $\mathbf{I}_{i}$ and $\mathbf{T}_{i}$ ($\forall i\in\{1,2,...,M\}$), respectively. The relevance of a page to the query is determined by calculating the inner product between the query vector and each modality’s representation: $z^{I}_{i}=\langle\mathbf{q},\mathbf{I}_{i}\rangle$ for the visual similarity and $z^{T}_{i}=\langle\mathbf{q},\mathbf{T}_{i}\rangle$ for the textual similarity. A straightforward method to calculate the final score $s_{i}$ is to combine these two scores through a weighted linear combination:
$s_{i}=\alpha z^{T}_{i}+(1-\alpha)z^{I}_{i},\quad\forall i\in\{1,2,...,M\},$ (4)
where $\alpha\in[0,1]$ is a weighting parameter that balances the contribution of the textual modality. This linear combination may yield sub-optimal results due to the inherent differences in modality nature and encoder capabilities. Also, the scales and distributions of the raw inner products are not directly comparable. To address this, we propose a unified co-modality–informed retrieval method that normalizes the scores into a shared, comparable space.
As shown in Fig. 2(c), to shift the inner product scores of visual and textual modalities to a comparable space, we first apply a sigmoid function to normalize the inner product scores to a [0,1] range:
$\bar{z}^{I}_{i}=\frac{1}{1+\exp\{-z^{I}_{i}\}},\quad\bar{z}^{T}_{i}=\frac{1}{1+\exp\{-z^{T}_{i}\}}.$ (5)
Empirically, we observed that the resulting distributions of $\bar{z}^{I}_{i}$ and $\bar{z}^{T}_{i}$ across a corpus approximate a Gaussian distribution. Therefore, to mitigate the bias from differing distribution parameters, we apply a Z-score normalization:
$\tilde{z}^{I}_{i}=\frac{\bar{z}^{I}_{i}-\mu^{I}}{\sigma^{I}},\quad\tilde{z}^{T}_{i}=\frac{\bar{z}^{T}_{i}-\mu^{T}}{\sigma^{T}},$ (6)
where $\mu^{I}=\frac{1}{M}\sum_{i=1}^{M}\bar{z}^{I}_{i}$ and $\sigma^{I}=\sqrt{\frac{1}{M}\sum_{i=1}^{M}(\bar{z}^{I}_{i}-\mu^{I})^{2}}$ are the mean and standard deviation of the visual similarity scores, respectively (with $\mu^{T}$ and $\sigma^{T}$ defined analogously for the text modality). This step ensures the scores from both modalities are standardized to a common scale with zero mean and unit variance. The final, unified retrieval score $\tilde{s}_{i}$ is then calculated as:
$\tilde{s}_{i}=\beta\tilde{z}^{T}_{i}+(1-\beta)\tilde{z}^{I}_{i},\quad\forall i\in\{1,2,...,M\},$ (7)
where $\beta$ represents the calibrated contribution of textual information. By eliminating the effects of distributional differences, the parameter $\beta$ more accurately denotes the relative confidence we place in the textual modality versus the visual modality for a given task.
3.3 Computational cost analysis
A critical requirement for a practical RAG system is low-latency retrieval. We emphasize that the CMRAG framework introduces negligible extra latency during the online retrieval phase, making it highly feasible for real-time applications. In a deployed system, all visual documents are parsed offline. In addition, all document images and texts are encoded offline by $\mathcal{E}^{I}$ and $\mathcal{E}^{T}$, respectively, and their embeddings are precomputed and stored in the retrieval index. At query time, the user’s query is encoded only once by $\mathcal{E}^{q}$. While the number of subsequent similarity calculations (inner products) is doubled compared to a single-modality retriever—as the query embedding is compared to both the text and image index—this operation is highly efficient on modern GPUs due to the parallelizable nature of matrix computations. Therefore, the online cost is dominated by a single encoding step, with the additional similarity calculations adding minimal overhead. This analysis confirms that CMRAG can achieve improved retrieval performance without a corresponding increase in computational cost.
4 Data Construction
To effectively train and evaluate our CMRAG pipeline on visual documents, we construct both training and evaluation datasets. Overall, the training data is constructed based on existing visual document RAG tasks, while the evaluation data covers mainstream benchmarks in visual document question answering. We will release all our processed datasets, including both training and evaluation sets, to facilitate further research by the community.
Training data. Our training corpus is built upon the synthetic dataset from VisRAG (Yu et al., 2025), which comprises diverse PDF documents (e.g., textbooks, academic papers, manuals) and approximately 240k query-document pairs generated by GPT-4o. To fully exploit the multimodal nature of these documents, we re-process all document pages (around 40k) using Qwen2.5-VL-7B (Bai et al., 2025) for end-to-end document parsing. This step extracts a structured representation for each page, including the entire page image, segmented sub-figures, and OCR-based text stored in HTML format. This enriched dataset provides fine-grained, aligned multimodal supervision, which is crucial for training our model to perform effective cross-modal retrieval and reasoning. Detailed data sources and processing prompts are provided in Appendix E.
Evaluation data. To comprehensively evaluate our method, we adopt several widely used VDQA benchmarks that span diverse domains and task types. Specifically, we include MMLongBench (Ma et al., 2024b) [^1] https://huggingface.co/datasets/MMlongBench/MMLongBench-Doc, LongDocURL (Deng et al., 2024) [^2] https://huggingface.co/datasets/THUDM/LongDocURL, and REAL-MM-RAG (Wasserman et al., 2025) [^3] https://huggingface.co/collections/ibm-research. These datasets cover industrial documents, scientific papers, presentation slides, and long multi-page documents, providing a broad and challenging testbed for assessing the retrieval and generation capabilities of our model. Detailed statistics of the evaluation datasets are provided in Tab. 1.
MMLongBench REAL-$M$-RAG LongDocURL
Doc FinReport FinSlides TechReport TechSlides Filtered
#Queries 1082 853 1052 1294 1354 770
#Documents 135 19 65 17 61 134
#Images 6529 2687 2280 1674 1963 11034
Avg. Query Length 95.59 78.56 93.38 89.03 82.17 161.83
Avg. Queries per Doc 8.01 44.89 16.18 76.12 22.20 5.75
Avg. Images per Doc 48.36 141.42 35.08 98.47 32.18 82.34
Table 1: Statistics of evaluation datasets. We apply a unified parsing procedure to all datasets, and sample one-third of LongDocURL for testing to balance controllability and cost.
@@ 5.1 Experiment setup
Baselines. We evaluate the effectiveness of our proposed method against several strong embedding model baselines: (1) BGE: A state-of-the-art text embedding model that uses only the parsed textual content from the documents (Xiao et al., 2023); (2) CLIP-B/32: The CLIP model with a Vision Transformer-B/32 image encoder; (3) CLIP-L/14-336: A larger and higher-resolution CLIP variant with a Vision Transformer-L/14 image encoder pretrained on $\times336$ pixel images (Radford et al., 2021); (4) SigLIP: The SigLIP model, which forms the foundation of our architecture (Zhai et al., 2023); and (5) SigLIP2: The latest iteration of the SigLIP model (Tschannen et al., 2025).
Evaluation metrics. We evaluate our method from two perspectives: (1) Retrieval metrics. We adopt standard measures including Recall, nDCG, and MRR. Detailed results for each dataset and metric are reported in the Appendix F. (2) Generation metrics. Traditional evaluation measures such as Exact Match (EM) and token-level F1 often fail to capture the semantic equivalence of multi-span or diverse gold answers, thereby underestimating model performance. To overcome this limitation, we employ a large language model (LLM) as an automatic judge to assess whether a generated response is correct, providing a fairer and more reliable comparison across different models.
Implementation details. We implemented our CMRAG framework using the Qwen2.5 series of models. The backbone Vision Language Model (VLM) for both document parsing and final answer generation is Qwen2.5-VL-7B-Instruct (Bai et al., 2025). For evaluation, we employ Qwen2.5-7B-Instruct (Yang et al., 2025a) as the judge model. To ensure deterministic and reproducible outputs, we set the decoding temperature to 0.0 for all models. Our unified encoding model produces embeddings with a dimensionality of 1152, and the maximum length of the text encoder is 768. For retrieval, the top-$k$ value is set to 3. All models are implemented using the PyTorch framework and trained on NVIDIA H100 GPUs.